OpenStudy (anonymous):

jade hiked 1 1/4 mile on the trail. Gretta hiked 1 1/2 times as far on the trail as Jade. How far did Gretta hike on the trail?

OpenStudy (anonymous):

@tootzrll

OpenStudy (anonymous):

\[\large \sf 1\frac{1}{4}=\frac{5}{4}\]\[\large \sf 1\frac{1}{2}=\frac{3}{2}\] \[\large \sf \frac{5}{4} \times \frac{3}{2}=?\]

OpenStudy (anonymous):

15/8

OpenStudy (anonymous):

Correct. You can turn it into a mixed fraction if needed

OpenStudy (litzyceron):

oh yea yur right @LegendarySadist lol

OpenStudy (anonymous):

1 7/8

OpenStudy (anonymous):

is that the answer in a mixed number